OpenStudy (anonymous):

Use the distributive property to multiply. A. (2y+4)(3y^2+y-4) Simplify the product using the distributive property B. (x-2)(x+4)

OpenStudy (anonymous):

do you know what to do?

OpenStudy (anonymous):

kind of just keep making simple mistakes and now im kinda mad at it ha

OpenStudy (anonymous):

just take your time: 2y*3y^2=6y^3 2y*y=2y^2 2y*-4=-8y 4*3y^2=12y^2 4*y=4y 4*-4=-16 6y^3+2y^2-8y+12y^2+4y-16= 6y^3+14y^2-4y-16

OpenStudy (anonymous):

so im still lost @~Jane_Doe~

OpenStudy (whpalmer4):

\[(2y+4)(3y^2+y-4)\]applying the distributive property: \[(2y+4)(3y^2+y-4) = 2y(3y^2+y-4) + 4(3y^2+y-4)\]Is that step clear?

OpenStudy (anonymous):

yeah i get it

OpenStudy (whpalmer4):

So now we do that again, as many times as we need to to get rid of all of the ()'s. \[2y*3y^2 + 2y*y + 2y*(-4) + 4*3y^2 + 4*y + 4*(-4)\]\]What do you get when you wrestle that to the floor?

OpenStudy (anonymous):

2^2+10x-4x-40=x^2+6x-40

OpenStudy (anonymous):

wait im ways off i used the wrong last number

OpenStudy (whpalmer4):

no equals sign in the middle, either

OpenStudy (whpalmer4):

also, why are you typing x when the problem uses y?

OpenStudy (anonymous):

oh ok im a little lost and i dont know why i put x's

OpenStudy (whpalmer4):

Let's go through it step by step: \[2y*3y^2 + 2y*y + 2y*(-4) + 4*3y^2 + 4*y + 4*(-4)\] \[2y*3y^2=\]

OpenStudy (anonymous):

I got that im at the point before the answer 6y^3+14y^2-4y-16

OpenStudy (whpalmer4):

\[6y^3 + 2y^2 -8y + 12y^2 + 4y -16\]Is that where you are?

OpenStudy (anonymous):

yeah that looks correct what do i do next

OpenStudy (whpalmer4):

add all the \(y^3\) terms together, then all the \(y^2\) terms together, then all the \(y\) terms together, and then all the numeric terms, that's called collecting like terms.

OpenStudy (whpalmer4):

Sorry, I have to leave now...